# Stronger, Fewer, & Superior: Harnessing Vision Foundation Models for Domain Generalized Semantic Segmentation

### v4

@@ 4.2 Comparison with State-of-The-Art Methods
In this section, we comprehensively evaluate Rein over five datasets within three generalization settings: GTAV normal-$\rightarrow$ Citys + BDD + Map, GTAV + Synthia normal-$\rightarrow$ Citys + BDD + Map, and Citys normal-$\rightarrow$ BDD + Map. Rein is benchmarked against state-of-the-art (SOTA) methods, which can be classified into two groups, including domain generalized semantic segmentation (DGSS) methods[69, 4, 49, 8, 29, 50, 35, 61, 22, 47, 65, 12], and parameter-efficient fine-tuning (PEFT) approaches [23, 21, 5].
@@
Investigation and comparison of various VFMs. Our analysis of VFMs and proposed Rein in the GTAV normal-$\rightarrow$ Citys + BDD + Map setting is presented in Tables 1 and 2. In this setup, models are fine-tuned using GTAV and evaluated on Cityscapes, BDD100K, and Mapillary. Note that, due to the fixed and relatively small number of trainable parameters in the decode head (20.6M), the count of trainable parameters presented in the tables are focused solely on the backbone and the PEFT module. Our results, as detailed in Table 1, indicate that frozen VFMs significantly outperform previous DGSS methods without specialized design. Moreover, as shown in Table 2, VFMs with full parameter fine-tuning exhibit enhanced performance relative to their frozen counterparts. Remarkably, Rein achieves even superior generalization capabilities, surpassing the full parameter fine-tuning with merely an extra 1% of trainable parameters compared to the original backbone. Visual samples for qualitative comparison are given in Fig. 3.
@@
Comparing Rein with SOTA on identical backbones. We conduct a comprehensive performance comparison of the proposed Rein against existing DGSS and PEFT methods under the GTAV normal-$\rightarrow$ Citys + BDD + Map setting, as detailed in Table 3. Owing to the robust feature extraction capabilities inherent in VFMs, DGSS methods, which typically enhance generalizability through strong data augmentation or consistency constraints, (e.g., AdvStyle, PASTA, and GTR), do not exhibit significant performance improvement. On the other hand, PEFT methods have demonstrated notable advancements. For instance, AdaptFormer outperforms the “Freeze” baseline using EVA02 as the backbone, while VPT shows improved performance over “Full” with DINOv2. Employing the same backbones (DINOv2 and EVA02), proposed Rein achieves superior performance and surpass previous DGSS and PEFT methods.
Multi-source generalization. In this part, we compare Rein against other DGSS methods under GTAV + Synthia normal-$\rightarrow$ Citys + BDD + Map setting, in which networks are fine-tuned using both GTAV and Synthia datasets, and tested on Cityscapes, BDD100K, and Mapillary. As shown in Table 7, we report the performance of Rein employing two VFMs, EVA02 and DINOv2. Our results demonstrate that Rein significantly surpasses existing DGSS methods by a large margin in average mIoU (from $45.9\%$ to $65.2\%$).
@@
Cityscapes-to-other datasets generalization. The generalization from one real-world dataset to others is pivotal for practical applications in the field. To this end, we conduct experiments under the Citys normal-$\rightarrow$ BDD + Map generalization setting. In this context, Rein, when coupled with the DINOv2-Large, demonstrates superior performance across all datasets. This underscores the effectiveness of Rein in generalizing to diverse real-world scenarios.
@@ 4.3 Ablation Studies and Analysis
In this subsection, we conduct extensive ablation studies within two settings: GTAV normal-$\rightarrow$ Citys and GTAV normal-$\rightarrow$ Citys + BDD + Map. For all experiments, Rein is applied to two VFMs, i.e., EVA02 and DINOv2.
@@
Analysis of the key components. Table 9 is dedicated to thoroughly examining the effectiveness of each component within Rein, focusing on how they influence recognition performance across various semantic categories. In the GTAV normal-$\rightarrow$ Citys generalization setting, we sequentially incorporate different components of Rein and assess their impact on enhancing performance when applied to two VFMs, EVA02 and DINOv2. Interestingly, we observe that the “Freeze” baseline occasionally exhibit better recognition for specific categories, e.g., ‘road, sidewalk’, compared to the “Full” baseline. This suggests that VFMs lose some pre-training knowledge during fine-tuning, and “Freeze” helps to prevent. Similarly, our methods mitigate this knowledge forgetting. Furthermore, our methods show improved recognition capabilities for the majority of the 19 categories. For example, in recognizing ‘wall, motorcycle, bicycle’, our approach significantly outperforms both the “Full” and “Freeze” baselines.
@@
Overall, “Rein-core” boosts the average performance across 19 classes. Furthermore, “Rein-link”, as mentioned in Sec. 3.3, further boosts accuracy for certain objects, including ‘car, bus, train, motorcycle’, especially when DINOv2 serve as the backbone. The strategy of employing layer-shared MLP weights efficiently reduces the number of trainable parameters from 59.33M to 5.02M. Lastly, the incorporation of a low-rank token sequence not only further reduces the number of trainable parameters but also positively influences the performance of the model.
@@
Study on token length $m$. The core component of Rein is a set of learnable tokens $T\in\mathbb{R}^{m\times c}$. We explored various lengths $m$ for the token sequence, ranging from 25 to 200. As demonstrated in Fig. 4, models with $m=100$ and $m=150$ both achieve a strong mIoU of 64.3% when utilizing DINOv2 as the backbone, and models with $m=100$ achieve the optimal mIoU of 63.6% when using EVA02. We ultimately selected $m=100$ as the most suitable parameter, which is consistently applied in subsequent experiments.
@@
Study on rank $r$ As shown in Table 10, we turn our attention to the effect of rank $r$ on model performance. When employing EVA02 as the backbone, the peak performance is achieved at $r=16$. Similarly, with DINOv2 as the backbone, the optimal results are observed at $r=16$ and $r=32$. Consequently, unlike LoRA [21], we opt for a comparatively higher value of $r=16$ for our model.
@@
Speed, memory, and storage. For practical applications, training speed, GPU memory usage, and model storage requirements are crucial. Lower training speeds and reduced GPU memory usage are beneficial for development of new methods and adaptation for new tasks. As shown in Table 11, compared to “Full” baseline, proposed Rein improves training speed and reduces GPU memory usage. Additionally, Rein marginally increases the storage needs by only 0.01GB. A significant advantage of Rein is that models trained under different settings can share the same backbone parameters. This means that for deployment in diverse tasks and settings, we can only swap the rein weights (0.01GB) and head weights (0.08GB), rather than all parameters.
@@
Source mIoU
GTAV 66.4
+Synthia 68.1
+UrbanSyn 78.4
+1/16 of
Cityscapes train set
82.5
Table 12: Results on Cityscapes validation set.
5 Conclusions
In this paper, we assess and harness Vision Foundation Models (VFMs) in the context of DGSS. Driven by the motivation that Leveraging Stronger pre-trained models and Fewer trainable parameters for Superior generalizability, we first investigate the performance of VFMs under diverse DGSS settings. Subsequently, we introduce a robust fine-tuning approach, namely Rein, to parameter-efficiently harness VFMs for DGSS. With a fewer extra trainable parameters, Rein significantly enhance generalization ability of VFMs, outperforming SOTA methods by a large margin. Rein can be seamlessly integrated as a plug-and-play adapter for existing VFMs based on plain vision transformer architecture, improving generalization while making training efficient. Extensive experiments across various settings demonstrate the substantial potential of VFMs in the DGSS field, validating the effectiveness of proposed Rein in parameter-efficiently harnessing VFMs for DGSS.
[…]
Supplementary Material
Figure 5: The curves of training loss and test metrics display consistent trends across different VFMs and decode heads: intuitively, as trainable parameters increase from $0.00M(Freeze)\rightarrow 2.53M(Rein)\rightarrow 304.24M(Full)$, the training loss monotonically decreases, indicating that a greater number of trainable parameters indeed better fit the training dataset. However, the test metrics on the target dataset initially rise and then fall, forming an inverted U-shape. This pattern suggests that the “Full” baseline overfits the training data, leading to diminished test performance. These findings are aligned with our motivation that Leveraging Stronger pre-trained models and Fewer trainable parameters for Superior generalizability. The blue bar charts in the figure represent the average mIoU tested on the Cityscapes, BDD100K, and Mapillary datasets, while the yellow line denotes the training loss during fine-tuning on GTAV dataset.
6 Discussion about Fewer Trainable Parameters
Classical neural network theory [16, 17] points out that as model capacity increases, the empirical risk (or training risk) monotonically decreases, indicating an improved fit to training data. Conversely, the true risk (or test risk) typically exhibits a “U-shaped” curve, initially decreasing and then increasing, a phenomenon known as overfitting. From a modern viewpoint, the scaling law [27] suggests that on a smaller fixed dataset, performance stops to improve as model parameters increase, leading to overfitting.
Fig. 5 showcases a consistent trend across four different configurations. As trainable parameters increase from $0.00M(Freeze)\rightarrow 2.53M(Rein)\rightarrow 304.24M(Full)$, the training loss monotonically decreases. However, the test metrics on the target dataset peak with Rein, which employs 2.53 million parameters and incurs a sub-optimal training loss. In contrast, the “Full” baseline, despite recording the lowest training loss, only achieves sub-optimal test performance, a clear indicator of overfitting when compared to other setups. This observation aligns with the conclusions in  [16, 27], supporting ours observation that leveraging Stronger pre-trained models and Fewer trainable parameters can lead to Superior generalizability.
7 Other decode head
Our experiments on Rein employ the Mask2Former [7] decode head, which shares structures or core concepts with numerous methods in dense prediction tasks [44, 67, 37, 2, 6]. The universality of Mask2Former highlights the significance of our findings for a range of segmentation tasks, including instance and panoptic segmentation. Furthermore, to demonstrate Rein’s effectiveness in enhancing backbone generalization and its robustness across various decode heads, we conduct supplementary experiments using the popular SemFPN decode head [32], in the GTAVnormal-$\rightarrow$ Cityscapes + BDD100K + Mapillary setting.
As shown in Table 13, Rein surpasses the “Full” and “Freeze” baselines, employing 2.53 million trainable parameters within the backbone, while the SemFPN decode head comprises 1.63 million parameters. Owing to the absence of object queries in SemFPN, the “linking tokens to instance” mechanism, described in Sec.3.3, is not utilized, resulting in a reduction of Rein’s trainable parameters from 2.99 million to 2.53 million. When compared to the complete Rein configuration using the Mask2Former, using SemFPN achieves sub-optimal performance, evident in the 64.3% mIoU reported in Table 2 and 62.1% mIoU in Table 9, both implemented with DINOv2-Large. These findings guide our decision to focus on experiments involving Mask2Former in the main paper.
Backbone Fine-tune Method Trainable $*{}^{*}$ mIoU
@@ 7 Other decode head
Table 13: Performance Comparison with the proposed Rein with SemFPN [32] as Backbones under the GTAV normal-$\rightarrow$ Cityscapes (Citys) + BDD100K (BDD) + Mapillary (Map) generalization setting. Models are fine-tuned on GTAV and tested on Cityscapes, BDD100K and Mapillary. The best results are highlighted. *** denotes trainable parameters in backbones.

### v5

@@ 4.2 Comparison with State-of-The-Art Methods
In this section, we comprehensively evaluate Rein over five datasets within three generalization settings: GTAV $\rightarrow$ Citys + BDD + Map, GTAV + Synthia $\rightarrow$ Citys + BDD + Map, and Citys $\rightarrow$ BDD + Map. Rein is benchmarked against state-of-the-art (SOTA) methods, which can be classified into two groups, including domain generalized semantic segmentation (DGSS) methods[83, 5, 61, 10, 38, 62, 45, 74, 30, 59, 79, 14], and parameter-efficient fine-tuning (PEFT) approaches [31, 29, 7].
@@
Investigation of various VFMs. Our analysis of VFMs and proposed Rein in the GTAV $\rightarrow$ Citys + BDD + Map setting is presented in Tables 1 and 2. In this setup, models are fine-tuned using GTAV and evaluated on Cityscapes, BDD100K, and Mapillary. Note that, due to the fixed and relatively small number of trainable parameters in the decode head (20.6M), the count of trainable parameters presented in the tables are focused solely on the backbone and the PEFT module. Our results, as detailed in Table 1, indicate that frozen VFMs significantly outperform previous DGSS methods without specialized design. Moreover, as shown in Table 2, VFMs with full parameter fine-tuning exhibit enhanced performance relative to their frozen counterparts. Remarkably, Rein achieves even superior generalization capabilities, surpassing the full parameter fine-tuning with merely an extra 1% of trainable parameters compared to the original backbone. Visual samples for qualitative comparison are given in Fig. 3.
@@
Comparing Rein with SOTA. We conduct a comprehensive performance comparison of the proposed Rein against existing DGSS and PEFT methods under the GTAV $\rightarrow$ Citys + BDD + Map setting, as detailed in Table 4. Owing to the robust feature extraction capabilities inherent in VFMs, DGSS methods, which typically enhance generalizability through strong data augmentation or consistency constraints, (e.g., AdvStyle, PASTA, and GTR), do not exhibit significant performance improvement. On the other hand, PEFT methods have demonstrated notable advancements. For instance, AdaptFormer outperforms the “Freeze" baseline using EVA02 as the backbone, while VPT shows improved performance over “Full" with DINOv2. Employing the same backbones (DINOv2 and EVA02), proposed Rein achieves superior performance and surpass previous DGSS and PEFT methods.
@@
Real-to-Real generalization of Rein. The generalization from one real-world dataset to others is pivotal for practical applications in the field. To this end, we conduct experiments under the Citys $\rightarrow$ ACDC, Citys $\rightarrow$ Cityscapes-C, and Citys $\rightarrow$ BDD + Map generalization setting. As shown in Table 3 and 7, Rein, when coupled with the DINOv2-Large, demonstrates superior performance across all datasets. This underscores the effectiveness of Rein in generalizing to diverse real-world scenarios.
Synthetic-to-real generalization of Rein. As Tab. 11 illustrates, trained on synthetic UrbanSyn+GTAV+Synthia datasets, Rein achieved a 78.4% mIoU on the Cityscapes validation set. Further improvement is possible with additional synthetic data and higher-quality images generated by diffusion models, like [1]. This result can also be a valuable pre-trained weight for data-efficient training, reaching an $82.5\%$ mIoU with 1/16 of Cityscapes training set. This is a significant performance for semi-supervised semantic segmentation.
More backbones. We extend our analysis to integrating Rein with Convolutional Networks, such as ResNet and ConvNeXt, and smaller scale architectures like DINOv2-S/B. As shown in Table 8, our findings reveal that Rein exhibits remarkable performance with diverse backbones.
@@ 4.3 Ablation Studies and Analysis
We conduct extensive ablation studies within two settings: GTAV $\rightarrow$ Citys and GTAV $\rightarrow$ Citys + BDD + Map.
@@
Analysis of the key components. Table 6 is dedicated to thoroughly examining the effectiveness of each component within Rein. In the GTAV $\rightarrow$ Citys generalization setting, we sequentially incorporate different components of Rein and assess their impact. Interestingly, we observe that the “Freeze" occasionally exhibit better recognition for specific categories, e.g., ‘road, sidewalk’, compared to the “Full". This suggests that VFMs lose some pre-training knowledge during fine-tuning, and “Freeze" helps to prevent. Similarly, our methods mitigate this knowledge forgetting. Furthermore, our methods show improved recognition capabilities for the majority of the 19 categories. For example, in recognizing ‘wall, motorcycle, bicycle’, our approach significantly outperforms both the “Full" and “Freeze" baselines.
@@
Overall, “Rein-core" boosts the average performance across 19 classes. Furthermore, “Rein-link" further boosts accuracy for certain objects, including ‘car, bus, train, motorcycle’, especially for DINOv2. Employing layer-shared MLP weights and low-rank token sequence efficiently reduces the number of trainable parameters and positively influences the performance of the model.
Methods Backbone Trainable Parameters mIoU
BDD Map Avg.
IBN [59] ResNet50 [26]   23.58M 48.6 57.0 52.8
DRPC [79] ResNet50 [26]   23.58M 49.9 56.3 53.1
GTR [61] ResNet50 [26]   23.58M 50.8 57.2 54.0
SAN-SAW [62] ResNet50 [26]   23.58M 53.0 59.8 56.4
WildNet [45] ResNet101 [26]   42.62M 50.9 58.8 54.9
HGFormer [14] Swin-L [52] 196.03M 61.5 72.1 66.8
Freeze EVA02-L [18]    0.00M 57.8 63.8 60.8
Rein (Ours) EVA02-L [18]    2.99M 64.1 69.5 66.8
Freeze DINOv2-L [58]    0.00M 63.4 69.7 66.7
Rein (Ours) DINOv2-L [58]    2.99M 65.0 72.3 68.7
Table 7: Performance Comparison of the Rein against other DGSS methods under Cityscapes $\rightarrow$ BDD100K (BDD) +Mapillary (Map) generalization. The best results are highlighted.
Avg. mIoU
ResNet
ResNet
ConvNeXt
DINOv2
DINOv2
Full 38.9 46.1 52.2 51.8 56.7
Ours 46.6 46.3 55.5 55.7 59.1
Table 8: Results for ConvNets and smaller backbones.
@@
Study on token length $m$. The core component of Rein is learnable tokens $T\in\mathbb{R}^{m\times c}$. We explored various lengths $m$ for the token sequence, ranging from 25 to 200. As demonstrated in Fig. 4, models with $m=100$ and $m=150$ both achieve a strong mIoU of 64.3%. We ultimately selected $m=100$ as the most suitable parameter.
@@
Study on rank $r$. As shown in Table 12, we turn attention to the effect of rank $r$ on model performance. With DINOv2 as the backbone, the optimal results are observed at $r=16$ and $r=32$. Consequently, unlike LoRA [29], we opt for a comparatively higher value of $r=16$ for our model.
@@
Speed, memory, and storage. For practical applications, training speed, GPU memory usage, and model storage requirements are crucial. As shown in Table 13, compared to “Full" baseline, proposed Rein improves training speed and reduces GPU memory usage. A significant advantage of Rein is that models trained under different settings can share the same backbone parameters. This means that for switch in diverse tasks and settings, we can only store and swap the rein weights (0.01GB) and head weights (0.08GB), rather than all parameters.
@@
Rank $r$ 4 8 16 32 64
Params 2.67M 2.77M 2.99M 3.42M 4.28M
DINOv2 (Large)  [58] Citys 65.8 66.1 66.4 66.1 66.4
BDD 60.2 60.3 60.4 60.7 61.0
Map 65.2 65.1 66.1 65.9 65.0
Avg. 63.7 63.9 64.3 64.3 64.1
Table 9: Ablation study on lora dim $r$.
VFMs Method Training Time GPU Memory Storage
DINOv2 (Large) Full 11.2 h 14.7 GB 1.22 GB
Rein   9.5 h 10.0 GB 1.23 GB
Table 10: Training Time, GPU Memory, and Storage.
5 Conclusions
In this paper, we assess and harness Vision Foundation Models (VFMs) in the context of DGSS. Driven by the motivation that Leveraging Stronger pre-trained models and Fewer trainable parameters for Superior generalizability, we first investigate the performance of VFMs under diverse DGSS settings. Subsequently, we introduce a robust fine-tuning approach, namely Rein, to parameter-efficiently harness VFMs for DGSS. With a fewer trainable parameters, Rein significantly enhance generalizability of VFMs, outperforming SOTA methods by a large margin. Rein can be seamlessly integrated as a plug-and-play adapter for existing VFMs, improving generalization with efficient training. Extensive experiments demonstrate the substantial potential of VFMs in the DGSS field, validating the effectiveness of proposed Rein in harnessing VFMs for DGSS.
6 Acknowledgements
This work was supported in part by the Anhui Provincial Key Research and Development Plan 202304a05020072, in part by the Postdoctoral Fellowship Program of CPSFGZB20230713, and in part by the National Natural Science Foundation of China under Grant 61727809.
[…]
Supplementary Material
Figure 5: The curves of training loss and test metrics display consistent trends across different VFMs and decode heads: intuitively, as trainable parameters increase from $0.00M(Freeze)\rightarrow 2.53M(Rein)\rightarrow 304.24M(Full)$, the training loss monotonically decreases, indicating that a greater number of trainable parameters indeed better fit the training dataset. However, the test metrics on the target dataset initially rise and then fall, forming an inverted U-shape. This pattern suggests that the “Full" baseline overfits the training data, leading to diminished test performance. These findings are aligned with our motivation that Leveraging Stronger pre-trained models and Fewer trainable parameters for Superior generalizability. The blue bar charts in the figure represent the average mIoU tested on the Cityscapes, BDD100K, and Mapillary datasets, while the yellow line denotes the training loss during fine-tuning on GTAV dataset.
7 Fewer Trainable Parameters
Classical neural network theory [21, 25] points out that as model capacity increases, the empirical risk (or training risk) monotonically decreases, indicating an improved fit to training data. Conversely, the true risk (or test risk) typically exhibits a “U-shaped" curve, initially decreasing and then increasing, a phenomenon known as overfitting. From a modern viewpoint, the scaling law [36] suggests that on a smaller fixed dataset, performance stops to improve as model parameters increase, leading to overfitting.
Fig. 5 showcases a consistent trend across four different configurations. As trainable parameters increase from $0.00M(Freeze)\rightarrow 2.53M(Rein)\rightarrow 304.24M(Full)$, the training loss monotonically decreases. However, the test metrics on the target dataset peak with Rein, which employs 2.53 million parameters and incurs a sub-optimal training loss. In contrast, the “Full" baseline, despite recording the lowest training loss, only achieves sub-optimal test performance, a clear indicator of overfitting when compared to other setups. This observation aligns with the conclusions in  [21, 36], supporting ours observation that leveraging Stronger pre-trained models and Fewer trainable parameters can lead to Superior generalizability.
Source Domain Cityscapes mIoU
GTAV 66.4
+Synthia 68.1
+UrbanSyn 78.4
+1/16 of Cityscapes Training set 82.5
Table 11: Results on Cityscapes validation set.
8 Value of synthetic data
As Tab. 11 illustrates, trained on synthetic UrbanSyn [24]+GTA+Synthia datasets, Rein achieved a 78.4% mIoU on the Cityscapes validation set. Further improvement is possible with additional synthetic data and higher-quality images generated by diffusion models, like [1]. This result can also be a valuable pre-trained weight for data-efficient training, reaching an $82.5\%$ mIoU with 1/16 of Cityscapes training set. This is a significant performance for semi-supervised semantic segmentation.
9 Ablation on decode head
[…]
As shown in Table 14, Rein surpasses the “Full" and “Freeze" baselines, employing 2.53 million trainable parameters within the backbone, while the SemFPN decode head comprises 1.63 million parameters. Owing to the absence of object queries in SemFPN, the “linking tokens to instance" mechanism, described in Sec.3.3, is not utilized, resulting in a reduction of Rein’s trainable parameters from 2.99 million to 2.53 million. When compared to the complete Rein configuration using the Mask2Former, using SemFPN achieves sub-optimal performance, evident in the 64.3% mIoU reported in Table 2 and 62.1% mIoU in Table 9, both implemented with DINOv2-Large. As shown in Table 15, the Mask2Former brings the 11.7% mIoU for ResNet101.These findings guide our decision to focus on experiments involving Mask2Former in the main paper.
10 Ablation on EVA02
@@
Study on rank $r$ As shown in Table 12, with EVA02 as the backbone, the optimal results are observed at $r=16$.
Speed, memory, and storage. As shown in Table 13, compared to “Full" baseline, proposed Rein improves training speed and reduces GPU memory usage.
Backbone Fine-tune Method Trainable Params mIoU
@@
Table 14: Performance Comparison with the proposed Rein with SemFPN [41] as Backbones under the GTAV $\rightarrow$ Cityscapes (Citys) + BDD100K (BDD) + Mapillary (Map) generalization setting. Models are fine-tuned on GTAV and tested on Cityscapes, BDD100K and Mapillary. The best results are highlighted. * denotes trainable parameters in backbones.
Backbone Decoder Tune mIoU
ResNet101 [45] DeeplabV3plus Full 34.4
ResNet101 Mask2Former Full 46.1
DINOv2 Mask2Former Full 61.7
DINOv2 Mask2Former Ours 64.3
[…]
11 Multi-source generalization.
In this part, we compare Rein against other DGSS methods under GTAV + Synthia $\rightarrow$ Citys + BDD + Map setting, in which networks are fine-tuned using both GTAV and Synthia datasets, and tested on Cityscapes, BDD100K, and Mapillary. As shown in Table 16, we report the performance of Rein employing two VFMs, EVA02 and DINOv2. Our results demonstrate that Rein significantly surpasses existing DGSS methods by a large margin in average mIoU (from $45.9\%$ to $65.2\%$).
Methods Publication mIoU
Citys BDD Map Avg.
RobustNet [10] CVPR 21 37.7 34.1 38.5 36.8
PintheMem [38] CVPR 22 44.5 38.1 42.7 41.8
SAN-SAW [62] CVPR 22 42.1 37.7 42.9 40.9
WildNet [45] CVPR 22 43.7 39.9 43.3 42.3
DIGA [74] CVPR 23 46.4 33.9 43.5 41.3
SPC [30] CVPR 23 46.4 43.2 48.2 45.9
EVA02 - Frozen [19, 18] arXiV 23 55.8 55.1 59.1 56.7
EVA02 + Rein - 63.5 60.7 63.9 62.7
DINOv2 - Frozen [58] arXiV 23 64.8 60.2 65.2 63.4
DINOv2 + Rein - 68.1 60.5 67.1 65.2
Table 16: Performance Comparison of the proposed Rein against other DGSS methods under GTAV + Synthia $\rightarrow$ Cityscapes (Citys) + BDD100K (BDD) +Mapillary (Map) generalization.